1.)A person can take either of two routes to work, through Matteson or Richton Park. Both take on average 35 minutes, and travel

2.)

The variance of a sample of data was 20. Find the standard deviation. (Give two decimal digits)
Answer #1

1) Numerator df = 8 - 1 = 7

Denominator df = 6 - 1 = 5

For 7 numerator df, 5 denominator df and 5% significance level:

Critical value = 4.88

2) Standard deviation = \sqrt{Variance} = \sqrt{20} = 4.47